OpenStudy (anonymous):

Write the nth term of the following sequence in terms of the first term of the sequence. 10, 20, 40, . . .

OpenStudy (anonymous):

This is a geometric progression with first term 10 and common ratio 2

OpenStudy (anonymous):

the nth term of GP is:: \[ar ^{n-1}\]

OpenStudy (anonymous):

@amberkels55

OpenStudy (imstuck):

The formula for finding the nth term in a geometric sequence is\[a _{n}=a _{1}r ^{n-1}\

OpenStudy (imstuck):

That's \[a _{n}=a _{1}r ^{n-1}\]

OpenStudy (imstuck):

No sorry its just \[a _{n}=10(2^{n-1})\]So to find the fifth term in the series, say you would fill in\[a _{5}=10(2^{5-1})\]or\[a _{5}=10(2^{4})\]or\[a _{5}=10(16)\]which is 160. See how that works?

OpenStudy (imstuck):

The nth term of the following sequence in terms of the first term is \[a _{n}=10(2^{n-1})\]That's your answer...that's what you would write.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@IMStuck Yes thank you!